OpenStudy (yessgeorgie8):

Need help on trigonometry/precalculus :( @satellite73

OpenStudy (johnweldon1993):

What's the question?

OpenStudy (anonymous):

yes?

OpenStudy (yessgeorgie8):

@satellite73 @johnweldon1993

OpenStudy (anonymous):

\[\sin(35)=\frac{b}{25}\\ b=25\sin(35)\] and a calculator

OpenStudy (anonymous):

others are similar

OpenStudy (yessgeorgie8):

i got -10.70456674

OpenStudy (yessgeorgie8):

@satellite73